Fairfax Paint stock is expected to pay a dividend of 3.76 dollars in 1 year and a dividend of 3.56 dollars in 2 years. The stock is currently priced at 100.99 dollars and is expected to be priced at 106.97 dollars in 1 year. What is the price of Fairfax Paint stock expected to be in 2 years? The stock’s dividend is paid annually and the next dividend is expected in 1 year.
Answer #1

Price of stock now=(Price of stock in 1 year+Dividend in 1 year)/(1+return)

=>return=(106.97+3.76)/100.99-1=9.64%

Price of stock in 1 year=(Price of stock in 2 years+Dividend in 2 years)/(1+return)

=>Price of stock in 2 years=106.97*(1+9.64%)-3.56=113.7267